E2-5 Calouiate incoime and invostmont balance alfocation of excess to andervalued assots Pop Company acquired a 30 percent interest in Son on January 1 for $500,000 cash. Assume the cost ot the investment eoualh the ti velue of Sans net assets: Pop asigned the $125.00D excess of tair alue over book value ot the interest acquired to the following assets: Inventories Building Goodwill $25,000 (sold in the current year) $50,000 (4 year remaining life at January 1) $50,000 Daring the year Son reported net income of $200,000 and paid $50,000 dividends. REQUIRED 1. Determine Pops income from Son for the year 2. Determine the December 31 balance of the Investment in Son account
Answer #1

Solution:

Net Income of son reported for the year = $200,000

Income of son for the purpose of share of income of Pop = $200,000 - Excess of fair value over book value of inventory sold during current year - Additional depreciation on building

= $200,000 - $25,000 - ($50,000/4) = $162,500

Pop's income from son for the year = $162,500*30% = $48,750

Balance of investment in son account on December 31 = $500,000 + share in income - Share of dividend

= $500,000 + $48,750 - ($50,000*30%) = $533,750